# Fast and accurate emulation of complex dynamic simulators Junoh Heo

### v1

@@ 4.1 Lotka-Volterra model
Figure 5: RMSE (first plot), CRPS (second plot) of emulation algorithms to $P$ with forcing input $C$ and computation time (third plot) across 100 repetitions for Lotka-Volterra model.
Figure 5 demonstrates the results for the case in which $C(t)$ is treated as a forcing input. As anticipated, the emulation performance improves due to the availability of a known input, and the computation time is halved with 0.20403 seconds per repetition for the exact algorithm and 48.03827 seconds for the MC-based algorithm. These results confirm that the exact method remains computationally efficient in the presence of a forcing input.
@@ 4.2 Lorenz system
where $X$, $Y$, and $Z$ represent the system’s state variables, and $a$, $b$, and $c$ are system parameters that govern the system’s dynamics. Using the classic values $a=-8/3$, $b=-10$, and $c=28$, originally employed by Lorenz to study convection rolls, the system exhibits chaotic behavior. We focus on a simulation over $t\in[0,25]$, and the initial state $\mathbf{x}(t_{0})=(X(t_{0})=-1,Y(t_{0})=-1,Z(t_{0})=-1)$. Initial designs are sampled from $[-10,10]^{3}$.
Figure 6: RMSE (first to third plots), CRPS (fourth to sixth plots) of emulation algorithms to $P$ and $C$ and computation time (seventh plots) across 100 repetitions for Lorenz system.
Figure 7: Lorenz system (black solid line) and emulation results of prediction (green dashed line) via each algorithm.
Figure 8: 3-d illustrations of Lorenz system (black solid line) and emulation results of prediction (green dashed line) via each algorithm.
From Figure 6, the emulation results show slight differences between the two algorithms. This discrepancy arises because the Lorenz system is highly sensitive to small changes, which can lead to significant differences, a phenomenon known as the “butterfly effect”. As shown, the exact algorithm can emulate the system efficiently, requiring only 0.95777 seconds per repetition, whereas the MC approach takes significantly longer at 143.22403 seconds. Figures 7 and 8 illustrate the predicted trajectories for each state variable, $X$, $Y$, and $Z$. Both algorithms provide accurate predictions at early time steps up to $t=20$, but their performance deteriorates beyond this point. For highly sensitive dynamic systems such as the Lorenz system, the exact algorithm is particularly advantageous, as MC methods introduce variations that exacerbate prediction errors over time. This underscores the importance of exact emulation for improving accuracy and computational efficiency in chaotic systems.
@@ 5 Discussion
In this paper, we propose fast and exact algorithms for emulating complex dynamic simulators. These algorithms enable efficient emulation compared to traditional MC-based approaches, while retaining their advantages. Numerical studies demonstrate that our method achieves equivalent results with significantly reduced computation time, over 100 times faster. Additionally, our approach accurately computes the predictive posterior mean and variance, preventing the accumulation of small differences from sampling errors.
From a computational perspective, the addition of jitter for numerical stability–commonly used to avoid ill-conditioned covariance matrices–may introduce minor discrepancies that could accumulate over time. In dynamic systems that are highly sensitive to small changes, such as those exhibiting the butterfly effect, these numerical artifacts can significantly impact the emulation outcomes. Furthermore, computational efficiency can be further improved by implementing low-level programming languages for operations.

### v2

@@ 4.2 Duffing oscillator
Figure 4: RMSE (first and second panels), MAE (third and fourth panels) of emulation algorithms to $x$ and $v$ and computation time (fifth panel) across 100 repetitions for the Duffing oscillator.
Figure 4 illustrates the Duffing oscillator results, which exhibit patterns similar to those observed for the Lotka–Volterra model. The MC approach again underperforms relative to the exact method, although the performance gap narrows as the number of samples increases. With 10 MC samples, RMSE and MAE are approximately 24–25% larger than those of the exact method; with 100 samples, the difference drops to 5–6%, and with 1000 samples, to about 2.5%. Computational times show the same trend as before, with the exact algorithm averaging 0.52205 seconds per repetition compared to 0.47833, 1.69159, and 107.25105 seconds for the MC approach with 10, 100, and 1000 samples, respectively.
Figure 5: Duffing oscillator (black solid line) and emulation results of prediction (green dashed line) via each algorithm with detected change points (red dotted vertical line)
Again, the change point detection results for the Duffing oscillator in Figure 5 follow the same pattern observed for the Lotka–Volterra model. Both methods yield accurate predictions, with predictive uncertainties around $10^{-6}$ for the exact method and $10^{-4}$ for the MC method. As in the previous case, the exact method does not detect any change points, whereas the MC method identifies changes at $t=12.35$ for $x$ and $t=5.51$ for $v$.
@@ 4.3 Lorenz system
where $X$, $Y$, and $Z$ represent the system’s state variables, and $a$, $b$, and $c$ are system parameters that govern the system’s dynamics. Using the classic values a=8/3a=-8/3, b=10b=-10, and $c=28$, originally employed by Lorenz to study convection rolls, the system exhibits chaotic behavior. We focus on a simulation over $t\in[0,25]$, and the initial state $\mathbf{x}(t_{0})=(X(t_{0})=-1,Y(t_{0})=-1,Z(t_{0})=-1)$. Initial designs are sampled from $10,10]^{3}$.
Figure 6: RMSE (first to third panels), MAE (fourth to sixth panels) of emulation algorithms to $X$, $Y$ and $Z$ and computation time (seventh panel) across 100 repetitions for Lorenz system.
Figure 6 shows that the Lorenz system results differ notably from those of the previous two simulators. The Lorenz system is a more sensitive and complex dynamical system, where small changes can lead to large deviations, a phenomenon known as the “butterfly effect”. This greater sensitivity makes the relative differences between the exact method and MC approaches more pronounced at small sample sizes and more inconsistent across state variables. With 10 MC samples, RMSE and MAE are 6–17% larger than those of the exact method, with 100 samples, the difference decreases to 1.5–4%, and with 1000 samples, it ranges from 0% to 3.5%. Computational costs remain substantially different, with the exact algorithm requiring 0.99940 seconds per repetition compared to 0.65037, 2.69461, and 160.60290 seconds for the MC approach with 10, 100, and 1000 samples, respectively. These results highlight that the MC approach requires a sufficiently large number of samples to achieve performance comparable to the exact method, but this comes at a substantial increase in computational time.
Figure 7: Lorenz (black solid line) and emulation results of prediction (green dashed line) via each algorithm with detected change points (red dotted vertical line) and corresponding 3-D trajectory illustrations.
Figure 7 illustrates the predicted trajectories with change point detection for each state variable, $X$, $Y$, and $Z$. Both algorithms produce accurate predictions at early time steps up to $t=20$, but performance deteriorates thereafter. For the exact method, the first time step at which the absolute difference between the prediction and the true output exceeds 10 is $t=19.88$ for $X$, $t=20.69$ for $Y$, and $t=19.99$ for $Z$, which is later than the corresponding times for the MC method, which are $t=19.77$ for $X$, $t=20.47$ for $Y$, and $t=19.90$ for $Z$. For highly sensitive dynamic systems such as the Lorenz system, the exact algorithm is particularly advantageous, as MC methods introduce variations that exacerbate prediction errors over time. This is consistent with the change point detection results, where the MC approach identifies a change point at $t=0.12$ for $X$ which is near the very beginning of the time steps, and at $t=19.29$ for $Y$ and $t=19.27$ for $Z$. In contrast, exact method maintains smaller uncertainties and detects only a single change point at $t=0.3$ for $X$ caused by increased uncertainty at the peak behavior, while no change points are detected for $Y$ and $Z$ within the given time steps under the same penalty value in the change point detection analysis.
@@ 5 Discussion
In this paper, we propose fast and exact algorithms for emulating complex dynamic simulators. These algorithms offer substantial improvements over traditional MC-based approaches by retaining their key advantages while significantly reducing computational overhead. Specifically, the proposed method bypasses MC approximations by deriving closed-form expressions for the posterior predictive mean and variance. This not only yields more than 150-fold improvements in computational efficiency without compromising accuracy, but also prevents the accumulation of small differences from sampling errors, thereby ensuring stable propagation over time. The framework also naturally accommodates forcing inputs by treating them as additional input variables within the GP formulation, enabling emulation of a broader class of dynamic systems. While uncertainty quantification is also supported in MC-based methods, our method preserves this capability and provides predictive variance analytically at each time step, supporting reliable uncertainty-aware emulation. To facilitate adoption and reproducibility, we provide an R package, dynemu, available on CRAN, which implements the proposed methodology in a user-friendly and computationally efficient framework.
From a computational perspective, the addition of jitter for numerical stability (Ranjan et al.,, 2011)–commonly used to avoid ill-conditioned covariance matrices–may introduce minor discrepancies that could accumulate over time. In dynamic systems that are highly sensitive to small changes, such as those exhibiting the butterfly effect, these numerical artifacts can significantly impact the emulation outcomes. Furthermore, computational efficiency can be further improved by implementing low-level programming languages for operations.